OpenStudy (anonymous):

Fun Land theme park has a total of 140 males and females who are employed at the park. If the number of males is 20 more than twice the number of females, how many males are employed at the park? i just need help setting up the equation

Parth (parthkohli):

\[m = males\\f = females\\\ \\ m + f = 140 \\ m = 20 + 2f \]

OpenStudy (anonymous):

let males = m and females = f m + f = 140 and m = 20 + 2*f find m

Parth (parthkohli):

And you could do that by isolating \(m\) in equation 1, then using substitution! :D

OpenStudy (anonymous):

m=100 f=40 ?

Parth (parthkohli):

You got it right again! Whee!

OpenStudy (anonymous):

yayyy thanks again, i have more that i need help with would you be able to help me set them up ?